OpenStudy (anonymous):

Simplify. Evaluate the numerical bases 3^2b^2c^-4 x 3^-1b^-5c^7

OpenStudy (ranga):

\[\Large 3^2b^2c^{-4} \times 3^{-1}b^{-5}c^7\]

OpenStudy (ranga):

Use the identity: \[\Large a^m \times a^n = a^{m+n}\]

OpenStudy (anonymous):

so what do you get when you simplify?

OpenStudy (ranga):

I will do one part and you do the rest so the next time you can do the whole thing yourself: 3^2 x 3^(-1) = 3^( 2 + (-1) ) = 3^1 = 3 What is b^2 x b^(-5) ? What is c^(-4) x c^7 ?

OpenStudy (anonymous):

\[\frac{9 b^2}{c^4} * \frac{c^7}{3 b^5}=\frac{3 c^3}{b^3} \]